OpenStudy (anonymous):

solve using quadratic formula x^2+3x+5 check answer please

OpenStudy (anonymous):

\[x=\frac{ -3\pm \sqrt{11} }{2}\]

OpenStudy (radar):

wouldn't that be a\[\sqrt{-11}\] rather than \[\sqrt{11}\]

OpenStudy (anonymous):

\[x=\frac{ -3\pm \sqrt{3} ^{2}-4*1*5}{2*1 } \]

OpenStudy (anonymous):

write -11

OpenStudy (radar):

Or \[\pm i \sqrt{11}\]

OpenStudy (radar):

The roots are complex.